I thought I submitted this one...but no answer yet.
Find the distance between (-3,-2) and (1,4).
If someone could show me how to do this, I would appreciate it!
Thank you!

Answer by checkley71(8403)

what you have is a right triangle with sides (4+2)=6 in the y direction
& (1+3)=4 in the x direction.
seeing as we're looking for the hypotenuse of this right triangle we use the pythagorean therom:a^2+b^2=c^2
6^2+4^2=c^2
36+16=c^2
c^2=52
c=sqrt52
c=7.21 is the distance between these 2 points.
------------------------------------------------
OR C=SQRT52=SQRT(4*13)=2SQRT(13)
THIS IS THE SAME ANSWER IN A DIFFERENT FORM.
YOU NEED TO BE ABLE TO CONVERT THESE TYPE OF ANSWERS AS YOU DIDN'T GIVE THE AVAILABLE ANSWERS.